ThanksGiving day for US people.

Installed new ff matrix ( acquired on 20111123_172201)

Calibrating time-history to not skip frames: fact=0.8 for a force max abs <= 0.62

System up @ 1305 UT ca. Verified time hist working with small skip frames.

start test 13:40 UT. End 13.55 ca. 10 times. All good.

Small jumps seen on actuators: 254, 513, 173 DSP. Logged in google DOCS

NOTE modified "recos.txt" changing disturb_len from 4000 to 4952 modified "idlctrl.conf" changing use_rpc from 0 to 1

PROCEDURE for TIMEHISTORY test:

* kill AOArb and AdsecArb

* set_plot, 'x' and test it with a plot, indgen(20). A plot should apply. if not, restart IDLCTRL, fsm_rip(), fsm_load_program()

* compile routines: .r test_time_hist2

* define turbulence file: histfile = '/home/aoacct/releases/src2/calib/adsec/672b/data/layer_phi_000000_time_hist_armando_672a.sav'

* set the shell: print, fsm_set_flat()

* start the turbulence test: print, test_time_hist2(hist=histfile)

* check stroke and forces into the log: tail /local/aolog/current/idlctrl.L.log. If ok answer OK to the dialog.

Reference values:
idlctrl.L          |INF|       907|2011-11-24 13:33:50.169664|             MAIN >  . Min,Max, RMS Force Hist  [N]: 0.092 0.114

idlctrl.L          |INF|       907|2011-11-24 13:33:50.169664|             MAIN >  . Min,Max, Pos  Hist  [um]: -4.3   4.9

idlctrl.L          |INF|       907|2011-11-24 13:33:50.169664|             MAIN >  . Min,Max Force Total [N]:-0.730 0.501

idlctrl.L          |INF|       907|2011-11-24 13:33:50.169664|             MAIN >  . Min,Max Pos  Total [um]: 42.7  69.9

* some skip will occurr. no problem.

* for get back the data: print, test_time_hist_get() and select the data_to_reduce.sav file into $ADOPT_MEAS/adsec_calib/disturbance/TN

(NOTE: in this case the time history will continue but without icc)

Time history test from 14:00 UT to 1520. One rest after downloaded data to reset back the ICC.

Slew test again: 15.20 UT to 15.34 UT (10 times). All ok.

2 noise spectrum taken for astigmatism oscillation: 20111124_153523, 20111124_153547

Low order bump seems to be still there even if not seen in noise spectrum (not checked.)

-- MarcoXompero - 24 Nov 2011
Topic revision: r3 - 01 Dec 2011, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback